# Strengthening Fake News Detection: Leveraging SVM and Sophisticated Text Vectorization Techniques. Defying BERT? thanks:

### v1

@@ Abstract
The rapid spread of misinformation, particularly through online platforms, underscores the urgent need for reliable detection systems. This study explores the utilization of machine learning and natural language processing, specifically Support Vector Machines (SVM) and BERT, to detect news that are fake. We employ three distinct text vectorization methods for SVM: Term Frequency Inverse Document Frequency (TF-IDF), Word2Vec, and Bag of Words (BoW), evaluating their effectiveness in distinguishing between genuine and fake news. Additionally, we compare these methods against the transformer large language model, BERT. Our comprehensive approach includes detailed preprocessing steps, rigorous model implementation, and thorough evaluation to determine the most effective techniques. The results demonstrate that while BERT achieves superior accuracy with 99.98% and an F1-score of 0.9998, the SVM model with a linear kernel and BoW vectorization also performs exceptionally well, achieving 99.81% accuracy and an F1-score of 0.9980. These findings highlight that, despite BERT’s superior performance, SVM models with BoW and TF-IDF vectorization methods come remarkably close, offering highly competitive performance with the advantage of lower computational requirements.
@@ IV-A Performance of SVM with TF-IDF
The TF-IDF vectorizing method performed attaching an accuracy of 99.52% and an F1 score of 0.9949, the SVM classifier running a linear kernel These findings, shown in Table I, show that TF-IDF is likewise a useful technique for text vectorizing in false news identification.
@@ IV-E Comparison with BERT Base Model
We matched the results of our SVM model with a BERT base model educated on the same fake news dataset in order to assess its performance even more. In three epochs, the BERT basic model attained outstanding performance with 99% accuracy and f1 score.
@@
The remarkable performance of the confusion matrix in textbfFigure 6 for the BERT base model in fake news identification reveals With only a few misclassifications, the matrix indicates an incredibly high number of accurate classifications for both actual and fake news events. With just 7 false positives and 19 false negatives, the model specifically successfully finds 4689 actual news items and 4265 erroneous news items. This great degree of accuracy emphasizes how strong and dependable the BERT base model is for separating real from fraudulent news. The low number of misclassifications suggests that the model is well-tuned and able to manage the complexity and subtleties in the dataset.
This analysis highlights generally how well SVM and BERT models address the problem of fake news detection. While the BERT model gives great accuracy and robustness at the expense of higher processing needs, the SVM model offers a lightweight and fast alternative. Future studies might investigate hybrid techniques combining the advantages of several models to improve detection accuracy and robustness even more.
@@ V Conclusion
The widespread problem of false news seriously jeopardizes the integrity of knowledge sharing and society’s stability. This research has explored the efficacy of Support Vector Machines (SVM) coupled with various text vectorization methods—namely Bag of Words (BoW), Term Frequency-Inverse Document Frequency (TF-IDF), and Word2Vec—in the detection of fake news. Our comprehensive experiments demonstrate that SVM classifiers, particularly those utilizing BoW and TF-IDF, deliver outstanding performance in distinguishing fake news from genuine news. The SVM classifier with a linear kernel and BoW vectorization achieved the highest accuracy of 99.81% and an F1-score of 0.9980, illustrating its robustness and effectiveness. The TF-IDF vectorization method also performed exceptionally well, achieving an accuracy of 99.52% and an F1-score of 0.9949. Although the Word2Vec technique, which captures semantic relationships, showed slightly lower performance (96.54% accuracy and 0.9644 F1-score), it remains a valuable tool in the text analysis domain.
@@
Furthermore, this study compared the performance of SVM models with the BERT base model, which exhibited slightly superior performance, particularly in terms of precision and recall, achieving perfect scores across all metrics. Despite the BERT base model’s high computational requirements, its exceptional performance underscores the potential of advanced deep learning techniques in fake news detection.
Overall, this study highlights the efficacy of straightforward vectorization techniques such as BoW and TF-IDF when paired with SVM classifiers for fake news detection. The findings suggest that robust and efficient models can be developed using these methods, contributing significantly to the automated identification of false information in news articles. Future research could investigate hybrid models that integrate the strengths of SVM and deep learning approaches, as well as additional features such as user behavior and network analysis, to further enhance the accuracy and reliability of fake news detection systems. By advancing the methodologies for detecting fake news, this study aims to support efforts in maintaining the credibility of information and curbing the spread of misinformation in the digital age.

### v2

# Strengthening False Information Propagation Detection: Leveraging SVM and Sophisticated Text Vectorization Techniques in comparison to BERT thanks:

@@ Abstract
The rapid spread of misinformation, particularly through online platforms, underscores the urgent need for reliable detection systems. This study explores the utilization of machine learning and natural language processing, specifically Support Vector Machines (SVM) and BERT, to detect fake news. We employ three distinct text vectorization methods for SVM: Term Frequency Inverse Document Frequency (TF-IDF), Word2Vec, and Bag of Words (BoW), evaluating their effectiveness in distinguishing between genuine and fake news. Additionally, we compare these methods against the transformer large language model, BERT. Our comprehensive approach includes detailed preprocessing steps, rigorous model implementation, and thorough evaluation to determine the most effective techniques. The results demonstrate that while BERT achieves superior accuracy with 99.98% and an F1-score of 0.9998, the SVM model with a linear kernel and BoW vectorization also performs exceptionally well, achieving 99.81% accuracy and an F1-score of 0.9980. These findings highlight that, despite BERT’s superior performance, SVM models with BoW and TF-IDF vectorization methods come remarkably close, offering highly competitive performance with the advantage of lower computational requirements.
@@ IV-A Performance of SVM with TF-IDF
The TF-IDF vectorizing method performed an accuracy of 99.52% and an F1 score of 0.9949, the SVM classifier running a linear kernel These findings, shown in Table I, show that TF-IDF is likewise a useful technique for text vectorizing in false news identification.
@@ IV-E Comparison with BERT Base Model
We matched the results of our SVM model with a BERT base model trained on the same fake news dataset to further evaluate its performance. In just three epochs, the BERT base model achieved excellent results, reaching an accuracy and F1-score of 99.98%.
@@
The confusion matrix in Figure 6 demonstrates the outstanding classification ability of the BERT base model, with only 7 false positives and 19 false negatives. The model correctly identified 4689 real news instances and 4265 fake news instances, showcasing its strong performance and ability to capture complex patterns within the dataset.
Computational Comparison: The BERT base model was trained using an NVIDIA Tesla T4 GPU with 16 GB VRAM, and required approximately 1 hour 45 minutes to complete 3 epochs. In contrast, the SVM model was trained on a standard CPU (Intel Core i7, 16 GB RAM) and completed training in under 3 minutes. This significant difference in training time and hardware requirements highlights the lightweight nature and computational efficiency of the SVM model, making it more suitable for scenarios with limited resources or real-time constraints.
This analysis illustrates how both SVM and BERT are effective for fake news detection. While BERT provides superior accuracy and robustness, it does so at the cost of high computational resources. On the other hand, SVM offers a faster and more efficient solution with reasonable performance. Future work could explore hybrid approaches that leverage the strengths of both models to further improve detection accuracy and robustness.
@@ V Conclusion
The widespread problem of false news seriously jeopardizes the integrity of knowledge sharing and societal stability. This research has explored the efficacy of Support Vector Machines (SVM) coupled with various text vectorization methods—namely Bag of Words (BoW), Term Frequency-Inverse Document Frequency (TF-IDF), and Word2Vec—for the detection of fake news. Our comprehensive experiments demonstrate that SVM classifiers, particularly those utilizing BoW and TF-IDF, deliver outstanding performance in distinguishing fake news from genuine news. The SVM classifier with a linear kernel and BoW vectorization achieved the highest accuracy of 99.81% and an F1-score of 0.9980, illustrating its robustness and effectiveness. The TF-IDF vectorization method also performed exceptionally well, achieving an accuracy of 99.52% and an F1-score of 0.9949. Although the Word2Vec technique, which captures semantic relationships, showed slightly lower performance (96.54% accuracy and 0.9644 F1-score), it remains a valuable tool in the text analysis domain.
@@
Furthermore, this study compared the performance of SVM models with the BERT base model, which exhibited slightly superior performance, particularly in terms of precision and recall, achieving near-perfect scores across all metrics. Despite the BERT base model’s high computational requirements, its exceptional performance underscores the potential of advanced deep learning techniques in fake news detection.
Overall, this study highlights the efficacy of straightforward vectorization techniques such as BoW and TF-IDF when paired with SVM classifiers for fake news detection. The findings suggest that robust and efficient models can be developed using these methods, contributing significantly to the automated identification of false information in news articles.
As a concrete direction for future work, hybrid approaches can be explored—such as leveraging BERT to generate contextual embeddings, followed by classification using lightweight models like SVM or logistic regression. This would combine the representational power of transformer models with the computational efficiency of classical machine learning. Additionally, integrating external signals like user behavior, propagation networks, or source credibility scores may further improve detection robustness. By advancing the methodologies for detecting fake news, this study aims to support efforts in maintaining the credibility of information and curbing the spread of misinformation in the digital age.